Show that for two scalar fields f and g:
∇.[∇f ×( f ∇g)]= 0.
1
Expert's answer
2019-04-09T10:36:08-0400
"\vec{∇}*[\vec{∇}f\times f\vec{∇}g]=[\vec{∇}\times\vec{∇}f] *f\vec{∇}g=0"

As

"[\vec{∇}\times\vec{∇}f]=curl(grad(f))=0"

We also used the property of triple product:


"\vec{a}*[\vec{b}\times\vec{c}]=[\vec{a}\times\vec{b}]*\vec{c}"